OpenStudy (susanna):

For all real numbers a and b, 2a • b = a2 + b2 Is this false or true? if its false explain.

OpenStudy (susanna):

@Janu16

OpenStudy (susanna):

hi

OpenStudy (susanna):

@AloneS

OpenStudy (evoker):

Definately false I would recommend just show a counter example by picking an a and b and checking

OpenStudy (susanna):

Why is it false

OpenStudy (evoker):

well first let me just check to make sure what you have is it this \[2ab=a^2+b^2\]

OpenStudy (susanna):

yes

OpenStudy (evoker):

well the first is two times a times b, the second is a times a plus b times b.

OpenStudy (susanna):

yup

OpenStudy (evoker):

So for instance if a is 3 and b is 4, the first is 2*3*4 or 24 while the second is 9+16 or 25

OpenStudy (susanna):

yea

OpenStudy (susanna):

so if put any numbers in for a and b it would be wrong?

OpenStudy (evoker):

well not any but most numbers.

OpenStudy (evoker):

for instance 1 and 1 works

OpenStudy (susanna):

ok that means that this equation is wrong because ifor example if you put a as 4 and b as 6 it would be wrong cause 2* 4*6 it would be 48 and the second would be 16+36 or 52

OpenStudy (evoker):

Yes, while it works for some cases it does not work for all. You have proved by contradiction that the statement is false.

OpenStudy (susanna):

thanks!!

OpenStudy (loser66):

In general: \(a^2+b^2 =2ab\\a^2-2ab+b^2=(a-b)^2=0\\\Longleftrightarrow a=b\) Therefore, for any \(a\neq b\) \(a^2+b^2 \neq 2ab\)